Suppose x is a real number such that 5cos(ϕ)=x where ϕ is a first quadrant angle. Find algebraic expressions (in terms of 5 and x ) for sin(ϕ) and tan(ϕ).
Answer by lwsshak3(11628)

Suppose x is a real number such that 5cos(ϕ)=x where ϕ is a first quadrant angle. Find algebraic expressions (in terms of 5 and x ) for sin(ϕ) and tan(ϕ).
***
5cosϕ=x
cosϕ=x/5
sinϕ=√(1-cos^2(ϕ))=√(1-x^2/25)=√((25-x^2)/25)=√(5^2-x^2)/5
tanϕ=sinϕ/cosϕ=√(5^2-x^2)/x